# Decentralized Coordination of Distributed Energy Resources through Local Energy Markets and Deep Reinforcement Learning

### v1

@@ Abstract
Transactive energy, implemented through local energy markets, has recently garnered attention as a promising solution to address the grid challenges in the form of decentralized, indirect demand response on a community level. Given the nature of these challenges, model-free control approaches, such as deep reinforcement learning, show promise for the decentralized automation of participation within this context. Existing studies at the intersection of transactive energy and model-free control primarily focus on socioeconomic and self-consumption metrics, overlooking the crucial goal of reducing community-level net load variability.
This study addresses this gap by training a set of deep reinforcement learning agents to automate end-user participation in ALEX, an economy-driven local energy market. In this setting, agents do not share information and only prioritize individual bill optimization. The study unveils a clear correlation between bill reduction and reduced net load variability in this setup. The impact on net load variability is assessed over various time horizons using metrics such as ramping rate, daily and monthly load factor, as well as daily average and total peak export and import on an open-source dataset. Agents are then benchmarked against several baselines, with their performance levels showing promising results, approaching those of a near-optimal dynamic programming benchmark.
@@
[inst2]organization=Computing Science, University of Alberta,city=Edmonton, postcode=T6G 2R3, state=AB, country=Canada
@@ 1 Introduction
To address these challenges, the research community has been actively exploring demand response (DR) methodologies. Broadly speaking, DR techniques leverage various signals to modulate end-user load demand, supporting electrical grid efficiency and reliability. These signals encompass both direct control commands to assets and incentive mechanisms intended to influence end-user behavior, thus delineating between direct and indirect DR. Notably, the key hurdles in indirect DR lie in aligning the interests of grid stakeholders and electricity end-users through appropriate incentive structures and subsequently ensuring sufficient participation to achieve the desired effect [4, 5, 6].
Traditionally, schedule-based approaches employing model predictive control (MPC) frameworks have been predominant in indirect DR. These approaches rely on behavioral models to form a forecast and then attempt to optimize load demand over a future time horizon. However, their inherent reliance on expert knowledge, high time complexity, and bias toward centralized information processing may impede their efficacy in addressing the rapid and disparate changes observed at the grid edge.
@@
This article addresses this research gap by training independent agents to automate end-user participation in LEMs and the utilization of DERs. The study demonstrates an emergent reduction in community net load variability even when agents solely prioritize individual bill optimization. To enhance benchmarking and future comparability, performance evaluation is conducted on an open-source dataset, and the agent’s performance is compared to several baselines. The trained DRL agents perform close to the near-optimal benchmark without information sharing or access to future information.
@@ 2.2 Autonomous Local Energy eXchange
Zhang et al. [35] delve into the essential properties required for ALEX’s settlement mechanism to incentivize RL agents to learn pricing in correlation with the settlement timestep $t_{\text{settle}}$ supply and demand ratios. Formulating ALEX as a mixed-form stochastic game suggests the existence of at least one Nash equilibrium. This insight facilitates the identification of a market mechanism possessing the desired properties through experiments that employ tabular Q-learning bandits under varied but fixed supply and demand ratios. Subsequent experiments deduce a market price function based on the current supply and demand ratio. A follow-up study by Zhang and Musilek [37] investigates a system incorporating a communal battery energy storage system (BESS) controlled by an expert-designed heuristic. The study demonstrates efficacy in avoiding violations of voltage-frequency constraints on a test circuit.
@@
2.3 Reinforcement Learning
Reinforcement Learning (RL) is a machine learning framework closely linked to optimal control paradigms.
As illustrated in Figure 1, RL focuses on optimizing the behavior of an agent that interacts with the environment through actions and subsequently receives observations and rewards.
@@ 2.3 Reinforcement Learning
This is typically formalized through the Markov Decision Process (MDP), represented by the tuple $(S,A,P_{a},R_{a})$. The MDP encapsulates the state space $S$, action space $A$, transition probabilities $P_{a}$ from state $s$ to the next state $s^{\prime}$ upon taking an action $a$, and receiving an immediate rewards $R_{a}$. A policy, denoted as $\pi$, characterizes an agent’s behavior through a probabilistic mapping from state $s$ to action $a$. For instance, this mapping could take the form of a Gaussian distribution, where the mean $\mu$ and standard deviation $\sigma$ are functions of the state $s$.
MDPs within the context of RL are typically time-discrete, allowing the notation of the time-step $t$ to represent a specific point in the interaction trajectory between the agent and the environment. This trajectory starts at $t=0$ and concludes at $t=T$. The return $G$ signifies the cumulative, discounted future reward,
@@
where $\gamma$ is the discount factor. The primary objective is to identify an optimal policy $\pi^{*}$ which maximizes the expected return $\mathbf{E}G$.
Distinguished from other MDP search methods by its emphasis on temporal difference and bootstrapping, RL agents iteratively learn the optimal policy $\pi^{*}$. They adjust their encoding in response to the reward signal received from the environment. The parameters underlying this encoding are denoted as $\theta$ and are updated through an RL learning algorithm’s loss function, often employing a stochastic gradient descent method. RL algorithms are generally categorized into two types: value-based and policy gradient methods. Value-based methods estimate state values $V$ or state-action values $Q$ and subsequently associate policies $\pi$ with these estimates. On the other hand, policy gradient methods directly learn policies $\pi$ or their parameters using a policy loss
$L(\theta)=\mathbb{E}\left[\log\pi_{\theta}(a_{t},s_{t})V_{t}\right],$ (4)
with actor-critic methods utilizing a critic to estimate state values $V$ and compute advantages $A$ in order to reduce variance, resulting in the corresponding actor-critic loss
$L(\theta)=\mathbb{E}\left[\log\pi_{\theta}(a_{t},s_{t})A_{t}\right],A_{t}=V_{t}-V_{\theta}(s_{t}).$ (5)
Deep Reinforcement Learning (DRL), an amalgamation of RL and deep neural networks, has gained traction for its ability to solve complex MDPs in a generalized manner [10, 11, 9]. DRL methods leverage replay buffers to store agent-environment interactions, facilitating multiple mini-batch stochastic gradient descent epochs. This necessitates the differentiation between the parameter set used to collect samples into the replay buffer $\theta_{old}$ and the new parameters $\theta$, which emerge as a result of gradient updates.
Particularly noteworthy within the dynamic landscape of DRL is Proximal Policy Optimization (PPO), introduced by Schulman et al. [40]. In contrast to naive actor-critic approaches, PPO employs a clipped surrogate objective based on the probability ratio $r(\theta)$. This ratio compares the probabilities of the new policy $\pi_{\theta}$ and the old policy $\pi_{\theta_{old}}$, aiming to mitigate policy drift and ensure the reliability of data collected into the replay buffer. PPO’s actor loss clips the magnitude of the policy ratio $r(\theta)$ within a tolerance parameter $\epsilon$
@@
In addition, most Proximal Policy Optimization (PPO) implementations incorporate generalized advantage estimation, a technique proposed by Schulman et al. [41], to reduce the variance of the advantage $A$.
@@ 3.1 Autonomous Local Energy eXchange as Markov Decision Process
However, future information, such as net load at settlement time $E^{b}_{t_{\text{settle}}}$, is not included in this observation space.
In contrast to the action space proposed by Zhang et al. [35], the action space $A^{b}$ for an agent at timestep $t$ exclusively includes the continuous battery action, scheduled for the future settlement time step $a_{BESS,t_{\text{settle}}}$. This action is constrained by the battery’s charge and discharge rates. The determination of bid and ask quantities at settlement time $t_{\text{settle}}$ relies on the residual net load, while bid and ask market conditions dictate prices following the round’s closure, guided by the price curve defined by Zhang et al. [35].
The building’s battery action $a^{b}_{BESS,t{\text{settle}}}$ is defined as a superposition of two components: the self-sufficiency maximizing, greedy battery action $a^{\pi_{0}}_{BESS,t_{\text{settle}}}$ and the agent’s learned action $a^{\pi_{\theta}}_{BESS,t_{\text{settle}}}$. Here, the policy $\pi_{0}$ represents the self-sufficiency maximizing policy, which aims to greedily minimize the amplitude of the participant’s net load $E^{b}_{t}$ using the residential BESS.
$a^{b}_{BESS,t_{\text{settle}}}:=a^{\pi_{0}}_{BESS,t_{\text{settle}}}+a^{\pi_{\theta}}_{BESS,t_{\text{settle}}}.$ (8)
This action and agent policy definition offers several distinct advantages, significantly expediting the learning process of the studied DRL agents. The policy $\pi_{0}$ can be computed at settlement time and serves as a reasonable initial heuristic, even though it may be far from the optimal policy. This approach enables more efficient state exploration while mitigating some of the internal environment modeling that the agent has to perform.
As a result, the agent’s reward function is formulated as the difference between the electricity bill $\textit{bill}^{b}_{t}$ and the bill incurred by the self-sufficiency maximizing policy $\pi_{0}$, denoted as $\textit{bill}^{b,\pi_{0}}_{t{\text{settle}}}$. This approach, in contrast to using the naive participant electricity bill $\textit{bill}^{b}_{t}$ as a reward signal, offers a clearer indication of whether the RL agents are learning a useful policy
$r^{b}_{t}:=\textit{bill}^{b}_{t_{\text{settle}}}-\textit{bill}^{b,\pi_{0}}_{t_{\text{settle}}}$ (9)
@@ 3.2 Shared Experience Recurrent Proximal Policy Optimization
The initialization of the actor’s final layer is designed to ensure that the mean $\mu$ exhibits an expected value of 0. This is achieved by sampling the weights and biases of this layer from a uniform distribution between 0.001 and -0.001. In a similar vein, the policy’s standard deviation $\sigma$ is initialized very narrowly. This setup enables the agent to commence training based on trajectories collected near the self-sufficiency maximizing policy $\pi_{0}$. This strategy is grounded in the assumption that the optimal policy $\pi^{*}_{\theta}$ is much closer to the self-sufficiency policy $\pi^{0}$ than to a pure random policy. Large deviations from $\pi^{0}$ are considered highly situational, while smaller deviations are more common. From a task decomposition perspective, the RL agents learn how to load shift to maximize self-sufficiency, an internally focused task, and then proceed to learn how to leverage the market, an externally focused task. Hence, this practice aims to bias the agents to first learn how to load shift and then learn how to utilize the market. Both adjustments contribute to notable improvements in convergence for the studied task.
Hyperparameters are used in this study are provided in Appendix A, along with a brief discussion of the tuning and monitoring process.
@@ 3.3 Experimental Design
The DRL agents are trained and evaluated on the CityLearn2022 dataset [50]. The open-source nature of this dataset enables subsequent studies to directly benchmark against this contribution across a diverse range of DR applications. This dataset provides a year of hourly data for 17 smart community buildings, featuring time series of energy demand, photovoltaic generation, and BESS performance characteristics. For each building, one independently acting agent is trained as outlined in Subsection 3.2. Therefore, one episode is defined as a full trajectory over the dataset and lasts 8760 steps, while one run fully trains such a set of 17 agents. For evaluation, the parameter set $\theta$ with the best episodic communal return $G^{B}$ is selected from a run, assuming that this snapshot represents the best-performing equilibrium between agents. This snapshot is updated throughout training when a new best communal return is achieved. From a set of 5 runs, the median performing run is selected for benchmarking purposes.
@@
The expectation is that ALEX RL shows a clear correlation between participant bill savings and improvement in the outlined performance metrics compared to the NoDERMS baseline. The desired outcome is for ALEX RL to perform comparably to ALEX DP. This achievement would indicate agent convergence to a near-optimal level of performance and a clear outperformance of the Individual DERMS benchmark. This outcome would effectively address the identified research gap by demonstrating a clear reduction in variability across the community due to participant automation DRL within a LEM. The achieved performance would be contextualized against a set of reasonable benchmarks.
@@ 4 Results and Discussion
This study aims to address a significant research gap highlighted in the background section by demonstrating a reduction in community-level variability of net load facilitated by DRL agents within a LEM. Towards this objective, this section establishes a clear connection between participant bill reduction and performance metrics within the chosen setting. Subsequently, a comparative analysis of the DRL agents against benchmarks introduced in the earlier subsection is conducted.
@@
In summary, the findings underscore the effectiveness of leveraging DRL agents in LEMs for load optimization. This emphasizes the potential for mitigating variability and optimizing energy consumption at a community level.

### v2

@@ Abstract
Transactive energy, implemented through local energy markets, has recently garnered attention as a promising solution to address the grid challenges in the form of decentralized, indirect demand response on a community level. Model-free control approaches, such as deep reinforcement learning (DRL), show promise for the decentralized automation of participation within this context. Existing studies at the intersection of transactive energy and model-free control primarily focus on socioeconomic and self-consumption metrics, overlooking the crucial goal of reducing community-level net load variability.
This study addresses this gap by training a set of deep reinforcement learning agents to automate end-user participation in an economy-driven, autonomous local energy market (ALEX). In this setting, agents do not share information and only prioritize individual bill optimization. The study unveils a clear correlation between bill reduction and reduced net load variability. The impact on net load variability is assessed over various time horizons using metrics such as ramping rate, daily and monthly load factor, as well as daily average and total peak export and import on an open-source dataset.
To examine the performance of the proposed DRL method, its agents are benchmarked against a near-optimal dynamic programming method, using a no-control scenario as the baseline. The dynamic programming benchmark reduces average daily import, export, and peak demand by 22.05%, 83.92%, and 24.09%, respectively. The RL agents demonstrate comparable or superior performance, with improvements of 21.93%, 84.46%, and 27.02% on these metrics. This demonstrates that DRL can be effectively employed for such tasks, as they are inherently scalable with near-optimal performance in decentralized grid management.
@@
[inst2]organization=Computing Science, University of Alberta & Alberta Machine Intelligence Institute,city=Edmonton, postcode=T6G 2R3, state=AB, country=Canada
@@ 1 Introduction
To address these challenges, the research community has been actively exploring demand response (DR) methodologies. Broadly speaking, DR techniques leverage various signals to modulate end-user net load demand, supporting electrical grid efficiency and reliability. These signals encompass both direct control commands to assets and incentive mechanisms intended to influence end-user behavior, thus delineating between direct and indirect DR. Notably, the key hurdles in indirect DR lie in aligning the interests of grid stakeholders and electricity end-users through appropriate incentive structures and subsequently ensuring sufficient participation to achieve the desired effect [4, 5, 6].
Traditionally, price schedule-based approaches have been predominant in indirect DR. Model predictive control (MPC) frameworks have been leveraged to determine such price schedules on a dynamic, state-dependent basis. These approaches rely on behavioral models to form a forecast and then attempt to optimize load demand over a future time horizon. However, their inherent reliance on expert knowledge, high time complexity, and bias toward centralized information processing may impede their efficacy in addressing the rapid and disparate changes observed at the grid edge.
@@
The main advantages of DRL are model independence, inherent adaptability, scalability, and the ability to learn online [13]. In general, these benefits must be weighted against the potential shortcomings of training duration, data requirements, and sample efficiency. However, these issues are not significant in the transactive market environment, which in itself provides a large amount of data stemming from continuous interactions of market participants.
@@
Building on prior studies by Zhang et al. [28] and May et al. [29], this article addresses this research gap by training independent agents to automate end-user participation in LEMs and the utilization of DERs. The main contributions of this study are as follows:
  • 1.
    Demonstrates a reduction in community net load variability, showing that selfish bill optimization by agents within LEMs can lead to emergent community-level benefits.
  • 2.
    Conducts performance evaluation on an open-source dataset, enhancing benchmarking and future comparability of the study’s results.
  • 3.
    Benchmarks DRL agents against several baselines, showing that the trained DRL agents achieve performance equivalent to a near-optimal dynamic programming benchmark.
@@ 2.2 Autonomous Local Energy eXchange
Zhang et al. [28] delve into the essential properties required for ALEX’s settlement mechanism to incentivize RL agents to learn pricing in correlation with the settlement timestep $t_{\text{settle}}$ supply and demand ratios. Formulating ALEX as a competitive mixed-form stochastic game suggests the existence of at least one Nash equilibrium. [^1] Note that this game is not zero sum — one agent does not succeed only at the expense of another. This insight facilitates the identification of a market mechanism possessing the desired properties through experiments that employ tabular Q-learning bandits under varied but fixed supply and demand ratios. Subsequent experiments deduce a market price function based on the current supply and demand ratio. A follow-up study by Zhang and Musilek [38] investigates a system incorporating a communal battery energy storage system (BESS) controlled by an expert-designed heuristic. The study demonstrates efficacy in avoiding violations of voltage-frequency constraints on a test circuit.
@@
2.3 Reinforcement Learning and Markov Decision Processes
Reinforcement Learning (RL) is a machine learning framework for optimizing agent interactions with an environment based on a reward signal, as illustrated in Figure 1.
@@ 2.3 Reinforcement Learning and Markov Decision Processes
This is formalized through the Markov Decision Process (MDP), represented by the tuple $(S,A,P_{a},R_{a})$. Here, $S$ denotes the state space, $A$ the action space, $P_{a}$ the transition probabilities from state $s$ to the next state $s^{\prime}$ upon taking an action $a$ and $R_{a}$ is the immediately received reward. A policy $\pi$ characterizes an agent’s behavior through a probabilistic mapping from state $s$ to action $a$, for example in the form of a Gaussian distribution, where the mean $\mu$ and standard deviation $\sigma$ are functions of the state $s$.
RL agents typically operate within a time-discrete MDP, where the time step $t$ represents points along the interaction trajectory, starting at $t=0$ and concluding at $t=T$. The return $G$ signifies the cumulative, discounted future reward,
@@
where $\gamma$ is the discount factor. The goal is to identify an optimal policy $\pi^{*}$ which maximizes the expected return $\mathbf{E}G$.
One of the key strengths of RL is its ability to learn optimal policies by directly interacting with the environment. This allows such agents to handle unknown transition dynamics and latent variables without requiring an explicit environment model. In contrast to other MDP search methods, RL agents iteratively learn the optimal policy $\pi^{*}$ through temporal difference learning and bootstrapping, adjusting their parameters $\theta$ in response to the reward signal. These parameters are updated via the RL algorithm’s loss function, typically using stochastic gradient descent. RL algorithms fall into two categories: value-based and policy gradient methods. Value-based methods estimate state values $V$ or state-action values $Q$, while policy gradient methods directly learn policies $\pi$ through a policy loss
$L(\theta)=\mathbb{E}\left[\log\pi_{\theta}(a_{t},s_{t})V_{t}\right].$ (4)
Actor-critic methods build upon policy gradients, reducing variance by relying on a critic to estimate state values $V$ and compute advantages $A$,
$L(\theta)=\mathbb{E}\left[\log\pi_{\theta}(a_{t},s_{t})A_{t}\right].A_{t}=V_{t}-V_{\theta}(s_{t}).$ (5)
Deep Reinforcement Learning (DRL), combining RL with deep neural networks, has become popular for solving complex MDPs [10, 11, 9]. DRL methods use replay buffers to store agent-environment interactions for multiple mini-batch gradient updates, requiring differentiation between the parameters used for sampling $\theta_{old}$ and the updated parameters $\theta$.
Proximal Policy Optimization (PPO), a popular actor-critic algorithm introduced by Schulman et al. [40], uses a clipped surrogate objective based on the probability ratio $r(\theta)$, comparing the new policy $\pi_{\theta}$ with the old policy $\pi_{\theta_{old}}$. This prevents policy drift and ensures reliable data collection. The actor loss clips the policy ratio $r(\theta)$ within a tolerance parameter $\epsilon$,
@@
Most PPO implementations also incorporate generalized advantage estimation (GAE) to reduce the variance of the advantage $A$, as proposed by Schulman et al. [41].
@@ 3.1 Autonomous Local Energy eXchange as Markov Decision Process
However, future information, such as net load at settlement time $E^{b}_{t_{\text{settle}}}$, is not included in this observation space. When designing this observation space, preliminary tests suggest the described feature set as a minimum necessary feature set that enables optimal decision-making.
In contrast to the action space proposed by Zhang et al. [28], the action space $A^{b}$ for an agent at timestep $t$ exclusively includes the continuous battery action, scheduled for the future settlement time step $t_{\text{settle}}$. This action is constrained by the battery’s charge and discharge rates. The determination of bid and ask quantities at settlement time $t_{\text{settle}}$ relies on the residual net load, while bid and ask market conditions dictate prices following the round’s closure, guided by the price curve defined by Zhang et al. [28].
The building policy $\pi_{b}$ is defined as a superposition of two policies:
  • 1.
    The self-sufficiency maximizing, greedy battery policy $\pi_{0}$ consisting only out of $\mu_{0}$.
  • 2.
    The agent’s learned policy $\pi_{\theta}$ with mean $\mu_{\theta}$ and $\sigma_{\theta}$.
resulting in a Gaussian distribution with
$\mu_{b}=\mu_{0}+\mu_{\theta},$ (8)
$\sigma_{b}=\sigma_{\theta}.$ (9)
This action and agent policy definition offers several distinct advantages, significantly expediting the learning process of the studied DRL agents. The policy $\pi_{0}$ can be computed at settlement time and serves as a reasonable initial heuristic, even though it may be far from the optimal policy. This approach enables more efficient state exploration while mitigating some of the internal environment modeling that the agent has to perform. The constraints to the agent’s actions are enforced by squashing $\pi_{b}$ into the appropriate range using $tanh$, similar to how SAC enforces a bounded policy [42].
The agents are tasked with minimizing their market-bound electricity bill $\textit{bill}^{b}$. The agent’s reward function is formulated as the inverse difference between the electricity bill $\textit{bill}^{b}_{t_{\text{settle}}}$ and the bill incurred by the self-sufficiency maximizing policy $\pi_{0}$, denoted as $\textit{bill}^{b,\pi_{0}}_{t{\text{settle}}}$,
$r^{b}_{t}:=\textit{bill}^{b,\pi_{0}}_{t_{\text{settle}}}-\textit{bill}^{b}_{t_{\text{settle}}}.$ (10)
Compared to using the naive participant electricity bill $\textit{bill}^{b}_{t}$ as a reward signal, this offers a clearer indication of whether the RL agents are learning a helpful policy since a positive reward value means the agents outperform the greedy policy baseline.
@@ 3.2 Shared Experience Recurrent Proximal Policy Optimization
In initial exploration, off-policy methods such as Soft Actor-Critic (SAC) [42] performed worse than on-policy methods such as PPO. The DRL algorihtm of choice for this study is therefore PPO, introduced in Section 2.3. It is a well-established algorithm with numerous existing and well-benchmarked implementations, which provide a reliable foundation for expansion and adaptation. Compared to value-based methods, PPO is better suited for the continuous action spaces present in the to-be-solved MDP. Furthermore, acting probabilistically can be beneficial in solving mixed-form stochastic games like ALEX. One of the key reasons for selecting RL over other applicable methods is its ability to handle uncertainty in the environment’s dynamics through model-free approaches.
@@
The initialization of the actor’s final layer is designed to ensure that the mean $\mu$ exhibits an expected value of 0. This is achieved by sampling the weights and biases of this layer from a uniform distribution between 0.001 and -0.001. In a similar vein, the policy’s standard deviation $\sigma$ is initialized very narrowly. This setup enables the agent to commence training based on trajectories collected near the self-sufficiency maximizing policy $\pi_{0}$. This strategy is grounded in the assumption that the optimal policy $\pi^{*}_{\theta}$ is much closer to the self-sufficiency policy $\pi^{0}$ than to a pure random policy. Large deviations from $\pi^{0}$ are considered highly situational, while smaller deviations are more common. From a task decomposition perspective, the RL agents learn how to load shift to maximize self-sufficiency, an internally focused task, and then proceed to learn how to leverage the market, an externally focused task. Hence, this practice aims to bias the agents to first learn how to load shift and then learn how to utilize the market. Both adjustments contribute to notable improvements in convergence for the studied task as identified in preliminary tests.
Additional auxiliary information such as the computational resources used, runtimes, hyperparameter values, and hyperparameter tuning, please refer to A.
@@ 3.3 Experimental Design
The DRL agents are trained and evaluated on the CityLearn2022 dataset [50]. The open-source nature of this dataset enables subsequent studies to directly benchmark against this contribution across a diverse range of DR applications. This dataset provides a year of hourly data for 17 smart community buildings, featuring time series of energy demand, photovoltaic generation, and BESS performance characteristics. As is common in studies like this, it is assumed that the larger grid operates as an infinite source and sink, able to absorb and supply all residual load without limitations at both the prosumer and community levels. For each building, one independently acting agent is trained as outlined in Subsection 3.2. Therefore, one episode is defined as a full trajectory over the dataset and lasts 8760 steps, while one run fully trains such a set of 17 agents. For evaluation, the parameter set $\theta$ with the best episodic communal return $G^{B}$ is selected from a run, assuming that this snapshot represents the best-performing equilibrium between agents. This snapshot is updated throughout training when a new best communal return is achieved. From a set of 5 runs, the median performing run is selected for benchmarking purposes.
@@
This experimental design aims to address the research gap identified in Section 2 by evaluating ALEX RL’s performance based on the outlined metrics. ALEX RL is compared against the NoDERMS baseline, with Individual DERMS serving as a reasonable performance benchmark, and ALEX DP representing a near-optimal solution.
@@ 4 Results and Discussion
This study aims to address a significant research gap highlighted in the background section by demonstrating a reduction in community-level variability of net load facilitated by DRL agents within a LEM. Towards this objective, this section establishes a clear connection between participant bill reduction and performance metrics within ALEX, assuming converged actors. Subsequently, a comparative analysis of the DRL agents against benchmarks introduced in the earlier subsection is conducted.
@@
In summary, the findings underscore the effectiveness of leveraging DRL agents in LEMs for load optimization. This emphasizes the potential for mitigating variability and optimizing energy consumption at a community level. The presented study narrowly focuses on evaluating the performance of a set of DRL agents within ALEX on a specific dataset. The established framework provides a decentralized, data-driven foundation that could potentially scale to larger, more diverse energy communities. However, the investigation of scaling this framework to broader applications may introduce additional challenges and, therefore, falls outside the scope of this work.